OpenStudy (anonymous):

the population of a city is 30,000 and increasing by 2.75% each year. when will the city reach 60,000?

OpenStudy (campbell_st):

ok... so this a growth model... looks something like the compound interest formula A = P(1+ r)^n so 60000 = 30000(1 + 0.0275)^n you'll need logs to find the number of years.... n